OpenStudy (anonymous):

4y + 1/9y^4 / y + 2/y^3

OpenStudy (rc94):

\[\frac{ 4y+1 }{ 9y ^{^{4}} }\times \frac{ y ^{3} }{ y+2 }\] I think this is equivalent, and then you can cancel down

OpenStudy (anonymous):

Thank s